Calculate [OH -] and pH for each of the following solutions.

(a) 0.0010 M NaOH

[OH-] = M pH =

(b) 0.0267 g of LiOH in 550.0 mL of solution

[OH -] = M pH =

(c) 69.4 mL of 0.00138 M Sr(OH)2 diluted to 600 mL

[OH -] = M pH =

(d) A solution formed by mixing 49.0 mL of 0.000590 M Sr(OH)2 with 77.0 mL of 4.2 x 10-3 M NaOH

[OH -] = M pH =
Answer #1

اه هم م ه ه . ) Naon M 10-3 Lo-) - (5-3 5- ) ر ا م log 10-3 per pon = // = 3 = ۱۶ = م/ 3=p خود , ره کی د هکر ادا 3، ۵ . 04 روe fordilution M, V. = More 69.480.00 138 = M2 x bow feyx0.00138 - D = 60 ots Ooools -)=2xo.ooo15 pon = -log[on- ] реч - пр. вplz give up thumb for rating.